# Research impact evaluation based on effective authorship contribution sensitivity: h-leadership index

### v1

@@ 4.2 Research collaboration trends
We find that the research output across Go8 is overwhelmingly multi-authored with 5 out of the 8 institutions having less than 4% single-authored papers and the remaining 3 being less than 7%. Despite the majority of publications by Go8 researchers having them placed as middle-authored, the top researchers at 7 of the 8 institutes published at least 42% of their output as single ( 3%), first ( 30%) or last ( 9%) authors, highlighting their significant contributions (Figure 6). UNSW leads the Go8 pack with the first and single-authored publications, followed by the University of Sydney(USyd). Surprisingly, the University of Melbourne (UniMelb) has shown to be the weakest when it comes to research leadership by first and single authors, although it ranked as the number 1 Australian university  [^8] https://about.unimelb.edu.au/facts-and-figures by QS  [^9] QS Ranking: https://www.topuniversities.com/university-rankings and Times of High Education (THE)  [^10] THE: https://www.timeshighereducation.com/world-university-rankings ranking for almost a decade.
Next we present an analysis of our h-leadership index for assessing the research impact of leading researcher profiles and compare with conventional metrics such as the h-index. We report Our findings quantitatively demonstrate the effectiveness h-leadership index (hl-index) in recognising the nuanced contributions of researchers within collaborative settings (Table 1) when compared to established indexes such as the h-index, hm-index, h-frac index and citations. Our hl-index has been able to effectively penalise the h-index as the median number of co-authors grows.
We find that the authorship distribution across Go8 is mostly consistent with the median number of coauthors  [^11] This is different from the statistics in the table. We computed the median number of coauthors in Table 1 by taking the average across each of the 50 author profiles whereas the box plot concatenates all 50 author publications before plotting. between 5 and 7 with Monash and the University of Sydney on the higher end (Figure 5).
@@ 4.3 Publication trends over time
We present the publication trend from January 1962 to December 2024 (Figure 7 and observe that the highest number of articles being published around 2018-2021 possibly indicating that Scopus does not cover recent year publications adequately or the top researchers have been reducing their publication output in recent years. We cannot determine the exact reason behind this behaviour. All universities also achieve a very high median number of citations ¿20 per publication with values ranging between 21 to 28 as shown in Figure 8.
@@ 4.4 Subject area distribution
We next present an analysis of the research impact for different subject areas of the top 50 cited researchers from Go8 Australian universities. We note that the median number of citations received by the researchers per paper is 25 (Figure 1) and the top 5 subject areas are: i) Medicine (98,785), ii) Biochemistry, Genetics and Molecular Biology (65,959), iii) Engineering (38,958), iv) Physics and Astronomy (33,289), v) Materials Sciences (24,792) (Table 3). The Go8 universities perform well-balanced research across all disciplines but when considering only the top 50 researchers the distribution gets skewed in favour of specific disciplines. The top two subject areas based on the number of publications are given as follows. ANU leads in Physics Astronomy and Engineering; Monash, UA, UniMelb, UQ, USyd and UWA lead in Medicine and Biochemistry, Genetics and Molecular Sciences; UNSW leads in Medicine and Neuroscience (Figure  9). We find that Decision Sciences, Arts and Humanities, Economics, Econometrics and Finances, Dentistry and Business, Management and Accounting have significantly lower publications than the other subject areas. This inconsistency in subject area distribution may impact the bibliometric analysis, explaining significant variation between universities.
@@ 5 Discussion
We observe that the majority of publications are from life sciences, physical sciences health sciences, and social science in comparison, had fewer publications. Figure 9 presented differences in citation trends and research impact, eg. the impact factor of journals in the areas of medicine and computer science is generally much higher than arts and humanities. In general social sciences, and humanities publications receive a lower number of citations and are not catered in the measurement of research impact [31, 32]. Aiello et al. [33] discussed strategies for better evaluation of the impact of social science and humanities research. This is why metrics such as Field Weight Citation Index (FWCI) [34] are becoming prominent with Scopus profiles providing this index for individual researchers and also institutions [35]. Therefore, an institution that is heavily producing research papers in areas of science and technology will naturally have a higher impact in terms of citations when compared to an institution specialising in arts and humanities. Considering this, the Indian government has created a separate ranking system based on institution type (i.e. Universities and Technical Institutions) , which needs to be considered for a better system. The ranking by THE and QS University Ranking [36] does not take into account the nature of the institution in terms of areas of focus, the size of the university, and the funding received. The purpose of university ranking systems has been questioned [37] as it is natural to find institutions that have high funding with more research focus leading to higher ranking when compared to institutions that are focused on education delivery. Despite its advantages, the h-leadership index faces notable challenges:
@@
    Sensitivity to the author listing order: The method assumes that authors were listed in order of their contribution with the first and last authors being the most significant. Although this is used in science and engineering, it may not hold across all disciplines where authorship position conventions may differ.
@@
    Scopus data limitations: Discrepancies in author/coauthor affiliations and incomplete coverage of recent publications impact the accuracy of computed metrics. We further observed in our study, that some author’s citation counts varied by an order of magnitude when cross-referenced with Google Scholar. The most recent years’ publications also show limited coverage (Figure 7) which may be on account of Scopus considering only peer-reviewed journals, books and conference proceedings as sources. Preprints, datasets, software libraries and many other digital formats which aren’t formally published but receive high citations thus get left out.
@@
The c-score offered by Stanford’s top 2 % scientists list is a way forward [16]; however, it is an elitist approach where we find bibliometrics information only for the top researchers worldwide. This is demotivating for early and middle-career researchers, even in leading universities as they fail to be listed since a fairly large number of citations in key papers are needed which takes time to gather. The hl-index is a way forward, however, its implementation as a web resource and a service that is easily available can change the academic field. The use of hl-index can then take place in university recruitment and promotions, which can eliminate the need to be mega-author publications in huge collaboration groups where there is little contribution.
Future work could explore the application of the h-leadership index across various disciplines and its potential integration with other metrics to further refine the evaluation of collaborative research efforts. Ultimately, the adoption of such advanced metrics will contribute to a more balanced and inclusive recognition of scholarly contributions in the academic community. Disciplines such as pure mathematics typically follow the author alphabetically, which is a limitation for the hl-index. In future works, the hl-index can be amended so that author profiles can collectively define their weightage in papers for those that are listed alphabetically. It would also be the case that a single author may be contributing to different fields with alphabetical author listing being used in some papers, while in others, contribution-based listing is used. An advanced version of the hl-index would take all these into account. Finally, the hl-index is just the beginning and a lot can be done for better analysis of research impact.
@@ 6 Conclusion
We presented the h-leadership index that provides a significant advancement in the evaluation of research performance, particularly in the context of multi-authored publications. We used a complementary Gaussian weighting scheme in the h-leadership index effectively acknowledges the varying levels of contribution among authors, offering a more nuanced and fair assessment than existing metrics. Our analysis of the top researchers across Go8 demonstrates the h-leadership index’s ability to more accurately reflect the contributions of both leading and supporting authors. This metric not only enhances the recognition of first and last authors but also ensures that middle authors are fairly considered, albeit with appropriate weight adjustments.

### v2

@@ 4.2 Research collaboration trends
We find that the research output across Go8 is overwhelmingly multi-authored, with 5 out of the 8 institutions having less than 4% single-authored papers and the remaining 3 being less than 7%. Despite the majority of publications by Go8 researchers having them placed as middle-authored, the top researchers at 7 of the 8 institutes published at least 42% of their output as single ( 3%), first ( 30%) or last ( 9%) authors, highlighting their significant contributions (Figure 6). UNSW leads the Go8 pack with the first and single-authored publications, followed by the University of Sydney(USyd). Surprisingly, the University of Melbourne (UniMelb) has shown to be the weakest when it comes to research leadership by first and single authors, although it ranked as the number 1 Australian university  [^8] https://about.unimelb.edu.au/facts-and-figures by QS  [^9] QS Ranking: https://www.topuniversities.com/university-rankings and Times of High Education (THE)  [^10] THE: https://www.timeshighereducation.com/world-university-rankings ranking for almost a decade.
Next, we present an analysis of our h-leadership index for assessing the research impact of leading researcher profiles and compare it with conventional metrics such as the h-index. We report Our findings quantitatively demonstrate the effectiveness of the h-leadership index (hl-index) in recognising the nuanced contributions of researchers within collaborative settings (Table 1) when compared to established indexes such as the h-index, hm-index, h-frac index and citations. Our hl-index has been able to effectively penalise the h-index as the median number of co-authors grows.
We find that the authorship distribution across Go8 is mostly consistent with the median number of coauthors  [^11] This is different from the statistics in the table. We computed the median number of coauthors in Table 1 by taking the average across each of the 50 author profiles, whereas the box plot concatenates all 50 author publications before plotting. between 5 and 7 with Monash and the University of Sydney on the higher end (Figure 5).
@@ 4.3 Publication trends over time
We present the publication trend from January 1962 to December 2024 (Figure 7 and observe that the highest number of articles was published around 2018-2021, possibly indicating that Scopus does not cover recent year publications adequately or that the top researchers have been reducing their publication output in recent years. We cannot determine the exact reason behind this behaviour. All universities also achieve a very high median number of citations greater than 20 per publication with values ranging between 21 to 28 as shown in Figure 8.
@@ 4.4 Subject area distribution
We next present an analysis of the research impact for different subject areas of the top 50 cited researchers from Go8 Australian universities. We note that the median number of citations received by the researchers per paper is 25 (Figure 1) and the top 5 subject areas are: i) Medicine (98,785), ii) Biochemistry, Genetics and Molecular Biology (65,959), iii) Engineering (38,958), iv) Physics and Astronomy (33,289), v) Materials Sciences (24,792) (Table 3). The Go8 universities perform well-balanced research across all disciplines, but when considering only the top 50 researchers, the distribution gets skewed in favour of specific disciplines. The top two subject areas based on the number of publications are given as follows. ANU leads in Physics, Astronomy and Engineering; Monash, UA, UniMelb, UQ, USyd and UWA lead in Medicine and Biochemistry, Genetics and Molecular Sciences; UNSW leads in Medicine and Neuroscience (Figure  9). We find that Decision Sciences, Arts and Humanities, Economics, Econometrics and Finances, Dentistry and Business, Management and Accounting have significantly fewer publications than the other subject areas. This inconsistency in subject area distribution may impact the bibliometric analysis, explaining significant variation between universities.
@@ 5 Discussion
We observe that the majority of publications are from life sciences, physical sciences, health sciences, and social science, which in comparison, had fewer publications. Figure 9 presented differences in citation trends and research impact, eg, the impact factor of journals in the areas of medicine and computer science is generally much higher than arts and humanities. In general, social sciences and humanities publications receive a lower number of citations and are not catered in the measurement of research impact [31, 32]. Aiello et al. [33] discussed strategies for better evaluation of the impact of social science and humanities research. This is why metrics such as the Field Weight Citation Index (FWCI) [34] are becoming prominent with Scopus profiles providing this index for individual researchers and also institutions [35]. Therefore, an institution that is heavily producing research papers in areas of science and technology will naturally have a higher impact in terms of citations when compared to an institution specialising in arts and humanities. Considering this, the Indian government has created a separate ranking system based on institution type (i.e. Universities and Technical Institutions) , which needs to be considered for a better system. The ranking by THE and QS University Ranking [36] does not take into account the nature of the institution in terms of areas of focus, the size of the university, and the funding received. The purpose of university ranking systems has been questioned [37] as it is natural to find institutions that have high funding with more research focus leading to higher ranking when compared to institutions that are focused on education delivery. Despite its advantages, the h-leadership index faces notable challenges:
@@
    Sensitivity to the author listing order: The method assumes that authors were listed in order of their contribution, with the first and last authors being the most significant. Although this is used in science and engineering, it may not hold across all disciplines where authorship position conventions may differ.
@@
    Scopus data limitations: Discrepancies in author/coauthor affiliations and incomplete coverage of recent publications impact the accuracy of computed metrics. We further observed in our study that some authors’ citation counts varied by an order of magnitude when cross-referenced with Google Scholar. The most recent years’ publications also show limited coverage (Figure 7), which may be on account of Scopus considering only peer-reviewed journals, books and conference proceedings as sources. Preprints, datasets, software libraries and many other digital formats that aren’t formally published but receive high citations thus get left out.
@@
The c-score offered by Stanford’s top 2 % scientists list is a way forward [16]; however, it is an elitist approach where we find bibliometrics information only for the top researchers worldwide. This is demotivating for early and middle-career researchers, even in leading universities, as they fail to be listed since a fairly large number of citations in key papers are needed, which takes time to gather. The hl-index is a way forward, however, its implementation as a web resource and a service that is easily available can change the academic field. The use of the hl-index can then take place in university recruitment and promotions, which can eliminate the need to be mega-author publications in huge collaboration groups where there is little contribution.
Future work could explore the application of the h-leadership index across various disciplines and its potential integration with other metrics to further refine the evaluation of collaborative research efforts. Ultimately, the adoption of such advanced metrics will contribute to a more balanced and inclusive recognition of scholarly contributions in the academic community. Disciplines such as pure mathematics typically follow the author alphabetically, which is a limitation for the hl-index. In future works, the hl-index can be amended so that author profiles can collectively define their weightage in papers for those that are listed alphabetically. It would also be the case that a single author may be contributing to different fields with alphabetical author listing being used in some papers, while in others, contribution-based listing is used. An advanced version of the hl-index would take all these into account. Finally, the hl-index is just the beginning, and a lot can be done for better analysis of research impact.
@@ 6 Conclusion
We presented the h-leadership index that provides a significant advancement in the evaluation of research performance, particularly in the context of multi-authored publications. We used a complementary Gaussian weighting scheme in the h-leadership index, which effectively acknowledges the varying levels of contribution among authors, offering a more nuanced and fair assessment than existing metrics. Our analysis of the top researchers across Go8 demonstrates the h-leadership index’s ability to more accurately reflect the contributions of both leading and supporting authors. This metric not only enhances the recognition of first and last authors but also ensures that middle authors are fairly considered, albeit with appropriate weight adjustments.